how do I find the percent increase of these numbers from 100 to 101 and from 36 TO 100
How do I find the percent decrease of these numbers from 112 to 98 and from 73 to 70

Answer by checkley71(8403)

(101-100)/100=1/100 OR 1 % INCREASE.
(100-36)/36=64/36=1.77777 OR 177.78% INCREASE.
(112-98)/112=14/112=.125=12.5% DECREASE.
(73-70)/73=3/73=.041=4.1% DECREASE.